BLUEFACE123:

What is probability of picking an orange marble and flipping a heads? erere Question 3 options: 30% 40% 15% 25%

Elite:

How many sides does it have? I know it's asking for the probability of flipping a heads but it's not two-sided so it can't be 50% since it's not an option.

MAGA:

how many marbles are their?

MAGA:

I guess it depends by how many marbles their are. Theirs a 50/50 Chance 4 getting a heads but ya just like I said it depends how many Marbles their are

BLUEFACE123:

here

1 attachment
MAGA:

so their are 10 marbles total and theirs 2 blue marbles.

BLUEFACE123:

Question 3 options: which one is the answer 30% 40% 15% 25%

Elite:

3/10 are orange marbles, and you still have a 50/50 chance on the coin.

MAGA:

@elite wrote:
3/10 are orange marbles, and you still have a 50/50 chance on the coin.
I forgot he was trying 2 get orange XD. Guess I didn't read it as well as I needed 2

BLUEFACE123:

so is it 30% yall gotta exlain it right

BLUEFACE123:

explain

MAGA:

So 3/10= 30% because 10/10=100% 9/10=90% 8/10=80% ect. If u go all the way down too 3/10 it'll be 30%

BLUEFACE123:

ight TY!!!!!!!!!!!!!!!!!!!!!!!

BLUEFACE123:

ight TY!!!!!!!!!!!!!!!!!!!!!!!

BLUEFACE123:

ight TY!!!!!!!!!!!!!!!!!!!!!!!

MAGA:

and the coin flip is a 50/50 chance so it dosn't really change the result

surjithayer:

\[P=\frac{ 3 }{ 10 }\times \frac{ 1 }{ 2}=\frac{ 3 }{ 20 }=15 \%\]